# Topological transition in cyanobacteria: from motion to structure

### v1

@@ Abstract
Many active systems are capable of forming intriguing patterns at scales significantly larger than the size of their individual constituents. We recently introduced a model for gliding filamentous cyanobacteria, which form a reticulate pattern as an early-stage biofilm. Our observations reveal that the filaments’ bodies follow the track of the head, and are subject to curvature fluctuations, defining a new class of active matter: active spaghetti. By identifying the dominant aspects of the filament’s dynamics, our model achieves great computational efficiency, without coarse-graining past the scale of an individual filament. Here, we explore large-scale collective effects and rich dynamics of cyanobacteria colonies, while still retaining information about the individual constituents’ dynamics and their interactions. We characterize the system’s topological transition from an isotropic distribution to a state of large-scale reticulate patterns by quantifying both dynamical and structural observables. Although it is not a periodic structure, in the steady state the reticulate pattern possesses a well-defined length scale determined by the filaments’ Péclet number, which controls the relative importance of activity and curvature fluctuations.
@@ I Introduction
Living matter is fundamentally characterized by its ability to self-organize [1, 2, 3, 4]. This is visible at length scales spanning a staggering range of about nine orders of magnitude, from the few microns of the complex structures present inside a eukaryotic cell, such as its nucleus, the endoplasmic reticulum, or the length of a microtubule, [5, 6, 7] to the complex spatial organization in bacterial biofilms extending over millimeters [8, 9, 10], to the scale of flocks of birds [11, 12] or schools of fish [13], up to kilometer-long herds of wildebeest [14, 15]. Starting from the seminal work of Vicsek et al. [16], the self-organization of large collectives of simple objects has garnered considerable interest within the field of active matter [16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Active systems are defined by the injection of energy into individual elements, which is converted into directed motion [26]. This energy injection explicitly breaks time-reversal symmetry [27] and allows active matter systems to exhibit emergent properties not observable in equilibrium systems [20, 21].
@@
Cyanobacteria are one of the most important [55, 56, 57] and ancient [58, 59] groups of species to have evolved on Earth, with evidence of their presence in the fossil record dating back to at least two billion years ago [60, 61]. Their ancestors first evolved oxygenic photosynthesis [57], which changed the chemical composition of the atmosphere, thus paving the way for the evolution of animals. In green plants and eukaryotic algae photosynthesis is carried out in the plastids, organelles now recognized to be endosymbiotic cyanobacteria, one of the two major examples of endosymbiosis in the history of life [62] (the other is the case of mitochondria). In terms of sheer number of cells, cyanobacteria are also the most common type of life that has ever lived [63]. They are found in extreme environments like hot springs, deserts and the polar regions [64], but also in everyday streams and fountains. A particularly diverse prokaryotic phylum, species of cyanobacteria range from unicellular to multicellular filamentous and branching forms [61, 65].
@@
Here, we explore the dynamics, structure, and phase behavior of a system of active filaments, modeled after the behavior of cyanobacteria. We investigate the large-scale collective effects of cyanobacteria colonies, while still retaining information about the individual constituents’ dynamics and their interactions. Combining numerical simulations with theoretical arguments, we uncover a rich dynamics for this model of gliding filaments as their Péclet number $\mathrm{Pe}$, density, and length are varied. As it develops towards a statistically steady state, this model shows transient behaviors that closely resemble corresponding experiments with filamentous cyanobacteria. We also find that the filaments’ effective translational diffusivity grows with the Péclet number, $\mathrm{Pe}$. The nonequilibrium phase diagram of the system exhibits a transition from isotropic to reticulate patterns as $\mathrm{Pe}$ and the area coverage, or density, increase. Finally, inspired by polymer physics, we compute the parallel and perpendicular pair correlation function for the filaments. The features of these correlation functions carries information about the structure of the reticulate network and faithfully predicts its characteristic length scale.
@@
Figure 1: Dynamics of gliding motility. (a) Close-up micrograph of Oscillatoria lutea filaments, composed of a chains of individual cells. (b) A time lapse of three micrographs demonstrates filament motion, by which each filament’s body glides along the track laid down by its head. The rightmost image highlights this path-tracking dynamics by overlaying the three images using shadings of different colors. (c) Schematic depiction of the model filaments, as discretized into chains of connected beads. Beads are placed at a fixed distance $v_{i}\Delta t$ apart. Upon advancing time by an increment $\Delta t$, the last bead of a filament is removed and placed as its new head. Non-reciprocal interactions occur when a filament’s head (filament $i$) is within a distance $\sigma$ from any bead of another filament, $j$. The head experiences a nematic aligning interaction $\mathcal{F}(\theta_{i})$ while filament $j$ remains unaffected by this interaction.
II Results
@@ II.1 Model for gliding cyanobacteria
Figure 2: Time evolution of systems of filaments forming into reticulate structures. The top row shows observations of experiments performed with Oscilatoria lutea, where $v_{0}=3\,\mu$m/s and $\tau=8\,$min [70]. The lower panels show simulations of a system initialized with randomly oriented filaments at $t=0$, using a filament length of $L=1$, a Péclet number of $\mathrm{Pe}=3.5$ and an area coverage of $\Phi=1$. The field of view shown and time between snapshots match the experimental measurements of O. lutea (i.e. $t=15$ corresponds to 2 hours). In both cases the initial pattern forms with smaller features that coarsen slightly to achieve a stable length scale over a period of a few hours ($t\approx 50$).
Empirical observations guide the development of our model of the gliding motion of cyanobacteria, which then predicts the emergent behavior of cyanobacteria colonies. When gliding on surfaces, the head of a trichome lays out a path that the rest of the filament follows very closely; this is an analogue of the metameric motion observed in segmented animals, such as annelids and myriapods [52]. Figure 1(b) shows a time-series of micrographs of Oscillatoria lutea, which are then overlaid to demonstrate how the body of each filament follows the track of its head. To capture this behavior, we discretize the trichomes as chains of connected beads. In contrast to active polymer models, however, these beads do not have active forces applied tangent to the length of the filament, nor is there an independent random force (e.g. thermal noise) on each bead, nor any reaction force. Rather, the only forces relevant to the motion of a filament are those applied to its head. A filament moving at a constant gliding speed $v_{i}$ is then modeled as a string of beads with a fixed spacing $v_{i}\Delta t$, where $\Delta t$ is the time step of the numerical integration. To advance the filament by one time step, we remove the last bead in the chain (the tail), and add a new bead to the head. A similar approach has been taken to study the dynamics of isolated filaments [52], and of filaments on a lattice [76]. In this way, for each bead $\alpha$ on any filament $i$ the resulting path-tracking motion for the beads’ positions $\bm{x}_{i}^{\alpha}$ can be expressed as
$\bm{x}_{i}^{\alpha}(t)=\bm{x}_{i}^{\alpha-1}(t-\Delta t)\,\text{,}$ (1)
@@
$\frac{d\bm{x}_{i}}{dt}=v_{i}\bm{\hat{t}}_{i}.$ (2)
Here, $\bm{\hat{t}}_{i}\equiv(\cos\theta_{i},\sin\theta_{i})^{\mathrm{T}}$ is the unit vector tangent to the filament at the head’s position, and defines the head’s orientation $\theta_{i}$. The gliding speed $v_{i}$ of each filament is constant in time, and drawn from a normal distribution with average $v_{0}$ and standard deviation $\delta v$; this is in agreement with experimental observations [70], and avoids spurious synchronization effects.
The gliding motion of cyanobacteria is characterized by continuous functions for the orientation and curvature of the filaments. Because the curvature is mathematically well defined and empirically measurable, a simple Brownian process for the angle $\theta$ is not appropriate; rather, we use a modified Ornstein–Uhlenbeck process to capture the curvature diffusion of the filament paths [70]. In this case, the angular speed $\omega_{i}$ and orientation $\theta_{i}$ of the head of filament $i$ evolve as
@@
The last term in Eq. (4) describes the interactions, of strength $J$, between nearby filaments, where $\mathcal{F}(\theta_{i})=N_{ij}^{-1}\frac{\partial}{\partial\theta_{i}}\sum_{i\sim j}\cos(\theta_{i}-\theta_{j}^{\bar{\alpha}})$ is a Lebwohl–Lasher interaction inducing nematic alignment between the filaments, $N_{ij}$ is the number of filaments within an interaction distance $\sigma$ of the $i$-th filament’s head, $\sum_{i\sim j}$ denotes the sum over these filaments, and the orientation angle $\theta_{j}^{\bar{\alpha}}$ of the bead closest to the $i$-th head $\bm{x}_{i}$. The interaction has a short, finite range $\sigma$, which is consistent with the filaments’ effective diameter; when the distance between the head of filament $i$ and any bead of filament $j$ is less than $\sigma$, the orientation of the head of filament $i$ undergoes a deflection due to the Lebwohl–Lasher interaction, however, the bead of filament $j$ remains unperturbed. This makes the interactions fundamentally non-reciprocal, and it is predicated on the experimental observation that when two trichomes meet, only the incident filament’s head participates in the alignment process, while the other filament is not deflected [70]. Non-reciprocal interactions explicitly break detailed balance [82, 83] and are common in living active matter [84]. A schematic depiction of the non-reciprocal nematic interactions used in our model is shown in Fig. 1(c).
@@
Figure 3: Characterizing filament motion in simulations. (a) The mean square displacement (MSD) of the filaments’ heads exhibits clear ballistic and diffusive regimes, as shown here for different $\mathrm{Pe}$ at fixed $\Phi=1$ and $L=1$. The inset shows how the MSD converges to a long-time diffusive behavior, $\langle(\Delta x)^{2}\rangle\rightarrow 4D_{\text{eff}}\Delta t$. This implies that filaments meander throughout the network, rather than being trapped within specific looped paths. (b) The effective translational diffusion coefficient increases with Péclet number for the fully interacting system, and is consistent with the case of noninteracting filaments ($g=0$ in Eq. (7)); the dashed line represents the relation $D_{\mathrm{eff}}=\frac{1}{2}v_{0}^{2}\mathrm{Pe}^{2}$. (c) Filament length has no discernible effect on the MSD, as shown here for the case of $\mathrm{Pe}=3.5$.
Figure 4: The reticulate bundles are nematic in nature. The two panels show the same reticulate pattern, with filaments colored according to their local tangent (orientation) using (a) polar and (b) nematic mappings. For the polar case, no preferential direction is discernible within the filament bundles. In the nematic mapping, where orientations $\theta$ and $\theta+\pi$ are assigned the same color, the bundles appear in bright colors, highlighting their nematic structure.
To explore the general state diagram of our system, we nondimensionalize the equations of motion by setting $t=t_{0}\tilde{t}$, $\bm{x}=x_{0}\tilde{\bm{x}}$, and by choosing $t_{0}=\tau$, the autocorrelation time, as the characteristic unit of time, and $x_{0}=v_{0}\tau$ as the characteristic length. The nondimensional equations of motion read
$\frac{d\tilde{\bm{x}}}{d\tilde{t}}=$ $~{}\tilde{v}_{i}\bm{\hat{t}}_{i}$ (5)
@@
where $\tilde{v}_{i}=v_{i}/v_{0}$, the dimensionless interaction rate $g=J\tau$, and the Péclet number is defined as $\mathrm{Pe}=(v_{0}\tau\delta\kappa)^{-1}$. In our dimensionless units $\mathrm{Pe}$ coincides with $(\delta\kappa)^{-1}$. In the following, we will drop the tilde signs for ease of notation.

### v2

# Topological transition in filamentous cyanobacteria: from motion to structure

@@ Abstract
Many active systems are capable of forming intriguing patterns at scales significantly larger than the size of their individual constituents. Cyanobacteria are one of the most ancient and important phyla of organisms that has allowed the evolution of more complex life forms. Despite its importance, the role of motility on the pattern formation of their colonies is not understood. Here, we investigate the large-scale collective effects and rich dynamics of gliding filamentous cyanobacteria colonies, while still retaining information about the individual constituents’ dynamics and their interactions. We investigate both the colony’s transient and steady-state dynamics and find good agreement with experiments. We furthermore show that the Péclet number and aligning interaction strength govern the system’s topological transition from an isotropic distribution to a state of large-scale reticulate patterns. Although the system is topologically non-trivial, the parallel and perpendicular pair correlation functions provide structural information about the colony, and thus can be used to extract information about the early stages of biofilm formation. Finally, we find that the effects of the filaments’ length cannot be reduced to a system of interacting points. Our model proves to reproduce both cyanobacteria colonies and systems of biofilaments where curvature is transported by motility.
@@ I Introduction
Figure 1: Dynamics of gliding motility. (a) Close-up micrograph of Oscillatoria lutea filaments, composed of chains of individual cells. The scale bar is $30~{}\mu$m. (b) A time lapse of three micrographs demonstrates filament motion, by which each filament’s body glides along the track laid down by its head; the time lapse between each frame is 51 s, the scale bars in each frame are $100~{}\mu$m. The rightmost image highlights this path-tracking dynamics by overlaying the three images using shadings of different colors. (c) Schematic depiction of the model filaments, as discretized into chains of connected beads. Beads are placed at a fixed distance $v_{i}\delta t$ apart. Upon advancing time by an increment $\delta t$, the last bead of a filament is removed and placed as its new head. Non-reciprocal interactions occur when a filament’s head (filament $i$) is within a distance $\sigma$ (pink disk; not drawn to scale) from any bead of another filament, $j$. The head experiences a nematic aligning interaction $\mathcal{F}(\theta_{i})$ while filament $j$ remains unaffected by this interaction.
Living matter is fundamentally characterized by its ability to self-organize [1, 2, 3, 4]. This is visible at length scales spanning a staggering range of about nine orders of magnitude, from the few microns of the complex structures present inside a eukaryotic cell, such as its nucleus, the endoplasmic reticulum, or the length of a microtubule [5, 6, 7], to the complex spatial organization in bacterial biofilms extending over millimeters [8, 9, 10], to the scale of flocks of birds [11, 12] or schools of fish [13], up to kilometer-long herds of wildebeest [14, 15]. Starting from the seminal work of Vicsek et al. [16], the self-organization of large collectives of simple objects has garnered considerable interest within the field of active matter [16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Active systems are defined by the injection of energy into individual elements, which is converted into directed motion [26]. This energy injection explicitly breaks time-reversal symmetry [27] and allows active matter systems to exhibit emergent properties not observable in equilibrium systems [20, 21].
@@
Cyanobacteria are one of the most important [55, 56, 57] and ancient [58, 59, 60, 61] groups of species to have evolved on Earth, with evidence of their presence in the fossil record dating back to at least two billion years ago [62, 63]. Their ancestors first evolved oxygenic photosynthesis [57], which changed the chemical composition of the atmosphere, thus paving the way for the evolution of animals. In green plants and eukaryotic algae photosynthesis is carried out in the plastids, organelles now recognized to be endosymbiotic cyanobacteria, one of the two major examples of endosymbiosis in the history of life [64] (the other is the case of mitochondria). In terms of sheer number of cells, cyanobacteria are also the most common type of life that has ever lived [65]. They are found in extreme environments like hot springs, deserts and the polar regions [66], but also in everyday streams and fountains. A particularly diverse prokaryotic phylum, species of cyanobacteria range from unicellular to multicellular filamentous and branching forms [63, 67].
@@
Here, we explore the dynamics, structure, and phase behavior of a system of active filaments, modeled after the behavior of cyanobacteria. We investigate the large-scale collective effects of cyanobacteria colonies, while still retaining information about the individual constituents’ dynamics and their interactions. Combining numerical simulations with theoretical arguments, we uncover a rich dynamics for this model of gliding filaments as their Péclet number $\mathrm{Pe}$, area coverage $\Phi$, and filament length $L$ are varied. As it develops towards a statistically steady state, this model shows transient behaviors that closely resemble corresponding experiments with filamentous cyanobacteria. We also find that the filaments’ effective translational diffusivity grows with the Péclet number, $\mathrm{Pe}$. The nonequilibrium phase diagram of the system exhibits a transition from isotropic to reticulate patterns as Péclet number and area coverage, or density, increase. Finally, inspired by polymer physics, we compute the parallel and perpendicular pair correlation function for the filaments. The features of these correlation functions carries information about the structure of the reticulate network and faithfully predicts its characteristic length scale.
@@
II Methods
@@ II.1 Model for gliding cyanobacteria
Figure 2: Time evolution of systems of filaments forming into reticulate structures. (a)-(e) micrographs of experiments performed with Oscilatoria lutea, where $v_{0}=3\,\mu$m/s and $\tau=8\,$min [72]. (f)-(k) simulations of a colony initialized with randomly oriented filaments, using a filament length $L=1$, a Péclet number $\mathrm{Pe}=3.5$ and an area coverage $\Phi=1$. The field of view shown and time between snapshots match the experimental measurements of O. lutea (i.e. $t=2\,\mathrm{h}/\tau=15$ corresponds to 2 hours). The scale bars in the experimental images are 101010\,10mm and scale bars in simulations 101010\,10mm$/(v_{0}\tau)$. In both experiments and simulations the initial pattern forms with smaller features that coarsen slightly to achieve a stable length scale over a period of a few hours ($t\approx 50$).
Empirical observations guide the development of our model of the gliding motion of cyanobacteria, which then predicts the emergent behavior of cyanobacteria colonies. When gliding on surfaces, the head of a trichome lays out a path that the rest of the filament follows very closely; this is an analogue of the metameric motion observed in segmented animals, such as annelids and myriapods [52].
Figure 1(b) shows a time-series of micrographs of Oscillatoria lutea, which are then overlaid to demonstrate how the body of each filament follows the track of its head. To capture this behavior, we discretize the trichomes as chains of connected beads. In contrast to active polymer models, however, these beads do not have active forces applied tangent to the length of the filament, nor is there an independent random force (e.g. thermal noise) on each bead, nor any reaction force. Rather, the only forces relevant to the motion of a filament are those applied to its head. Filaments can glide over one another with no discernible friction, as such we do not include any steric repulsion in our quasi-2-dimensional (2D) model.
A filament moving at a constant gliding speed $v_{i}$ is modeled as a string of beads with a fixed spacing $v_{i}\delta t$, where $\delta t$ is the time step of the numerical integration. The gliding speed $v_{i}$ of each filament is constant in time, and drawn from a normal distribution with average $v_{0}$ and standard deviation $\delta v$; this is in agreement with experimental observations [72], and avoids spurious synchronization effects. Within any one simulation, all filaments have the same length $L$. Therefore, the number of beads in each filament is $L/(v_{i}\delta t)$. Because $v_{i}$ is normally distributed, different filaments will have a slightly different number of beads.
To advance the filament by one time step, we remove the last bead in the chain (the tail), and add a new bead as the new head. A similar approach has been taken to study the dynamics of isolated filaments [52], and of filaments on a lattice [78]. In this way, for each bead $\alpha$ on any filament $i$ the resulting path-tracking motion for the beads’ positions $\mathbf{x}_{i}^{\alpha}$ can be expressed as
$\mathbf{x}_{i}^{\alpha}(t)=\mathbf{x}_{i}^{\alpha-1}(t-\delta t)\,\text{,}$ (1)
@@
$\frac{d\mathbf{x}_{i}}{dt}=v_{i}\mathbf{\hat{t}}_{i}.$ (2)
Here, $\mathbf{\hat{t}}_{i}\equiv(\cos\theta_{i},\sin\theta_{i})^{\mathrm{T}}$ is the unit vector tangent to the filament at the head’s position, and defines the head’s orientation $\theta_{i}$.
The gliding motion of a cyanobacteria filament is characterized by its orientation and curvature. The curvature is mathematically well defined and has an empirically measurable exponential autocorrelation function with autocorrelation time $\tau$. A Wiener process acting on the angle $\theta$ does not satisfy these properties, that is, the curvature is not a well-defined quantity, and the autocorrelation of the curvature does not exist (and hence cannot be an exponential function). The simplest model that captures these properties is an Ornstein–Uhlenbeck process describing the curvature diffusion of the filament paths [72]. In this case, the angular speed $\omega_{i}$ and orientation $\theta_{i}$ of the head of filament $i$ evolve as
@@
Figure 3: Dynamics of filaments in simulations. (a) The mean square displacement (MSD) of the filaments’ heads exhibits clear ballistic and diffusive regimes, as shown here for different $\mathrm{Pe}$ at fixed $\Phi=1$ and $L=1$. The inset shows how the MSD converges to a long-time diffusive behavior, $\langle(\Delta x)^{2}\rangle\rightarrow 4D_{\text{eff}}\Delta t$. The orange crosses show the crossover time $t_{\times}=4D_{\text{eff}}/v_{0}^{2}$ between ballistic and diffusive regime. This implies that filaments meander throughout the network, rather than being trapped within specific looped paths. (b) The effective translational diffusion coefficient increases with Péclet number for the fully interacting system. Switching off the interactions ($g=0$ in Eq. (7)) increases slightly $D_{\mathrm{eff}}$, with the deviation growing with $\mathrm{Pe}$. The dashed line represents the relation $D_{\mathrm{eff}}=\frac{1}{2}v_{0}^{2}\tau\mathrm{Pe}^{2}$. (c) Filament length has no discernible effect on the MSD, as shown here for the case of $\mathrm{Pe}=3.5$.
Figure 4: The reticulate bundles are nematic in nature. The two panels show the same reticulate pattern of a simulation with $\mathrm{Pe}=3.5$, $\Phi=1$ and $L=1$ at steady state, with filaments colored according to their local tangent (orientation) using (a) polar and (b) nematic mappings. The scale bars are 101010\,10mm$/(v_{0}\tau)$. For the polar case, no preferential direction is discernible within the filament bundles. In the nematic mapping, where orientations $\theta$ and $\theta+\pi$ are assigned the same color, the bundles appear in bright colors, highlighting their nematic structure.
The last term in Eq. (4) describes the interactions, of strength $J$, between nearby filaments, where $\mathcal{F}(\theta_{i})=N_{ij}^{-1}\frac{\partial}{\partial\theta_{i}}\sum_{i\sim j}\cos(\theta_{i}-\theta_{j}^{\bar{\alpha}})$ is a Lebwohl–Lasher interaction [84, 85] inducing nematic alignment between the filaments, $N_{ij}$ is the number of filaments within an interaction distance $\sigma$ of the $i$-th filament’s head, $\sum_{i\sim j}$ denotes the sum over these filaments, and the orientation angle $\theta_{j}^{\bar{\alpha}}$ of the bead closest to the $i$-th head $\mathbf{x}_{i}$. The interaction has a short, finite range $\sigma$, which is consistent with the filaments’ effective diameter; when the distance between the head of filament $i$ and any bead of filament $j$ is less than $\sigma$, the orientation of the head of filament $i$ undergoes a deflection due to the Lebwohl–Lasher interaction, however, the bead of filament $j$ remains unperturbed. This makes the interactions fundamentally non-reciprocal, and it is predicated on the experimental observation that when two trichomes meet, only the incident filament’s head participates in the alignment process, while the other filament is not deflected [72]. Non-reciprocal interactions explicitly break detailed balance [86, 87] and are common in living active matter [88]. A schematic depiction of the non-reciprocal nematic interactions used in our model is shown in Fig. 1(c).
@@
Given the equation of motion (3) and (4), a Péclet number may be defined as $\mathrm{Pe}=v_{0}/(l\sqrt{D_{\omega}\tau})$ with a length scale $l$. This dimensionless number may be understood as a ratio of how quickly curvature $\kappa$ is transported within the system compared with the time it takes to randomize a filament’s curvature.
Figure 5: Emergence of reticulate patterns. The mosaic plot shows the steady-state behavior of the system at different area coverage $\Phi=0.25,0.5,0.75,1.0,1.25$ and Péclet numbers $\mathrm{Pe}=0.875,1.75,3.5,7.0,14.0$. The individual panels show simulation snapshots at steady state in a $30\times 30$ domain with periodic boundary conditions. Filament length is kept constant at $L=1$. The panels’ frames indicate the system’s average nematic order $\langle S\rangle$, calculated as an average of $1\times 1$ sized blocks. With increasing $\mathrm{Pe}$, the system transitions from a disordered state into one showing reticulate patterns of nematic bundles. The inset at $(\mathrm{Pe},\Phi)=(3.5,1.0)$ shows an example of the spatially resolved nematic order parameter $S_{\text{loc}}$; regions of higher density correlate with higher local nematic order parameter. Bundles show a high degree of ordering, whereas dilute regions in between are less ordered.
To explore the general state diagram of our system, we nondimensionalize the equations of motion by setting $t=t_{0}\tilde{t}$, $\mathbf{x}=l\tilde{\mathbf{x}}$, and by choosing $t_{0}=\tau$, the autocorrelation time, as the characteristic unit of time, and $l=v_{0}\tau$ as the characteristic length. The nondimensional equations of motion read
$\frac{d\tilde{\mathbf{x}}}{d\tilde{t}}=$ $~{}\tilde{v}_{i}\mathbf{\hat{t}}_{i}$ (5)
@@
where $\tilde{v}_{i}=v_{i}/v_{0}$, the dimensionless interaction rate $g=J\tau$, and the Péclet number naturally arises as $\mathrm{Pe}=1/\sqrt{D_{\omega}\tau^{3}}$. In dimensionless units the width of the curvature distribution is given by the inverse Péclet number $\mathrm{Pe}^{-1}$, such that filaments with high $\mathrm{Pe}$ are on average straighter. In the following, we will drop the tilde signs for ease of notation.
II.2 Simulations